As a Christmas Specialty, Lisette makes 120 pounds of candy to sell at 20.00per pound. she mixes candy which sells for 16.50 per pound with same amount of another candy which usually sells for 23.50 per pound. How many pounds of each will she usE?

Answer by mananth(16946)

percent ---------------- quantity
Candy type I 16.5 ---------------- x Lb
Candy type II 23.5 ---------------- 120 - x Lb
Mixture 20.00% ---------------- 120

16.5x+23.5(120- x)=20*120

16.5x+2820-23.5x=2400
16.5x-23.5x=2400-2820
-7x=-420
/-7
x=60 Lb $16.50 Candy type I
60 Lb $23.50 Candy type II